OpenStudy (anonymous):

while solving for the roots in y=x^5-5x using calculus i obtain imaginary roots what should i do? how can i graph if the roots are imaginary?

OpenStudy (anonymous):

any one could help me please?

OpenStudy (anonymous):

\[x^5-5x=0\]\[x(x^4-5)=0\implies x=0\]\[x^4=5\implies x=\sqrt[4]{5},\,-\sqrt[4]{5},\,i\sqrt[4]{5},\,-i\sqrt[4]{5}\]